OpenStudy (anonymous):

Gilberto needs to purchase baseballs and bats for his youth baseball league. He’ll need at least 15 pieces of equipment, but he only has $435 to spend. Baseballs cost $20 each and bats cost $35 each. Which is a combination of baseballs and bats that Gilberto can afford to buy? 9 baseballs and 7 bats 10 baseballs and 7 bats 8 baseballs and 6 bats 8 baseballs and 10 bats

OpenStudy (anonymous):

so if a = # of baseballs and b = # of bats you can set up two equations One: a + b >= 15 Two: 20a + 35b <= 435 now, you can do one of two things, you can solve the equations, or you can plug in the numbers they gave you into each equation and see which one fits. (I recommend the second suggestion)

OpenStudy (anonymous):

oh sorry i got this (: can you help me with another one?

OpenStudy (anonymous):

@tabby2014

OpenStudy (anonymous):

A group of 11 people are going to the movies one Saturday. Adult tickets cost $11.50 each and child tickets cost $7.25 each. If the total cost for admission to the movies is $105.25, how many children attended the movie? 4 5 6 7

OpenStudy (anonymous):

okay so a = # of adults and c = # of children a + c = 11 and 11.50a + 7.25c = 105.25 on this one, I would solve for a in the first equation and substitute that into the second equation and solve for c

OpenStudy (anonymous):

thanks @tabby2014 your awesome

OpenStudy (anonymous):

what should i do now @tabby2014

OpenStudy (anonymous):

for what?